A wooden block of mass 1.9kg rests on a rough horizontal surface. If the limiting frictional force is 12N, calculate the coefficient of friction, take g =10m/s².
1
Expert's answer
2020-11-23T10:26:30-0500

Applying the Newton’s Second Law of Motion, we get:


"\sum F_x=ma_x=0,""F_{appl}-F_{fr}=0,""F_{appl}=F_{fr}=\mu N = \mu mg,""\mu=\dfrac{F_{fr}}{mg}=\dfrac{12\ N}{1.9\ kg\cdot 10\ \dfrac{m}{s^2}}=0.63"

Answer:

"\mu=0.63"